# The gravitational-wave emission from the explosion of a 15 solar mass star with rotation and magnetic fields

### v1

@@ 1 Introduction
Core-collapse supernovae (CCSNe) are the explosive deaths of massive stars. Once the star’s iron core reaches its effective Chandrasekhar mass, the core collapses until it reaches and overshoots nuclear saturation density. Due to the stiffening to the nuclear equation of state (EoS), the collapse is then halted, and a shock wave is launched outwards due to the rebound of the iron core. Due to energy losses by nuclear dissociation and neutrino losses, the shock quickly stalls, but still expands as an accretion shock to a radius of $\sim 150$ km. Shock revival for the majority of CCSNe is expected to be powered by neutrinos, which need to be sufficiently strong to deliver explosion energies of $\sim 10^{51}$ erg as in observed CCSNe. Explosions powered by the neutrino-driven mechanism have been simulated extensively to understand the dynamics, remnant properties, neutrinos, and the gravitational-wave emission of CCSNe (see Müller (2020); Abdikamalov et al. (2020); Burrows & Vartanyan (2021) for recent reviews). However, the majority of 3D simulations of neutrino-driven explosions do not include the effects of rotation and magnetic fields.
@@
CCSNe are a promising multi-messenger source for the current ground-based gravitational-wave detectors Advanced LIGO (LIGO Scientific Collaboration et al., 2015), Advanced Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021). No gravitational waves have been detected from CCSNe in previous observing runs yet (e.g., Abbott et al., 2020; Szczepańczyk et al., 2023). Significant progress has been made in recent years predicting the gravitational-wave emission for neutrino-driven CCSNe (Powell & Müller, 2019; Burrows et al., 2020; Morozova et al., 2018; Andresen et al., 2017; Kuroda et al., 2017; Mezzacappa et al., 2020; Pan et al., 2021; Radice et al., 2019). Although simulations from different groups can differ in their predictions of the gravitational-wave amplitudes and precise frequency trajectories, all groups have found the dominant feature of the gravitational-wave signal to be the high frequency g/f-mode. Our understanding of the gravitational-wave emission modes from 3D simulations has enabled several groups to develop phenomenological models to aid CCSN gravitational-wave parameter estimation and searches (Astone et al., 2018; Powell & Müller, 2022; Bruel et al., 2023).
@@
The gravitational-wave signal from CCSNe with rotation and magnetic fields is not so well understood. Energetic explosions with rapid rotation may have enough gravitational-wave energy to be detected outside the Milky Way (Szczepańczyk et al., 2021). Rotation results in a spike in the time-series at the core-bounce time, which has been extensively studied (Dimmelmeier et al., 2002; Richers et al., 2017; Abdikamalov et al., 2014), including the early post-bounce phase in 3D (Scheidegger et al., 2008). However, gravitational-wave emission from the later phases of the explosion has not yet been investigated as thoroughly. Several studies have included rotation in their neutrino-driven explosions (Takiwaki & Kotake, 2018; Andresen et al., 2019; Powell & Müller, 2020; Shibagaki et al., 2020; Pan et al., 2021; Takiwaki et al., 2021), and a few gravitational-wave signals are now available for long-duration simulations that include both magnetic fields and rotation, where the magnetic fields played a major role in driving the explosion (Jardine et al., 2022; Bugli et al., 2023; Powell et al., 2023). The f/g-mode is also found to be the dominant feature in the gravitational-wave emission in simulations that include rotation and magnetic fields. However, as the effects of magnetic fields and rotation have so far been neglected in CCSN astroseismology studies, it is less clear how the gravitational-wave frequency relates to the proto-neutron star (PNS) properties. One noteworthy exception is the work of Raynaud et al. (2022), which addressed later phases beyond the first $\mathord{\sim}1\,\mathrm{s}$ of evolution, and identified a clear imprint of magnetic fields from an $\alpha-\Omega$-dynamo in the PNS convection zone on gravitational waves from inertial modes.
@@
The standing accretion shock instability (SASI) (Blondin et al., 2003; Blondin & Mezzacappa, 2006; Foglizzo et al., 2007) can also result in gravitational-wave emission in the most sensitive frequency band of current gravitational-wave detectors, and is often present in waveforms from neutrino-driven explosions. The gravitational-wave amplitude of the SASI-driven modes is usually lower, and more difficult to detect, than the dominant g/f-mode. However, the majority of previous works investigating SASI signatures in the gravitational-wave signal have not included both magnetic fields and rotation. CCSN simulations without neutrino transport have shown that SASI-driven turbulence can amplify magnetic fields exponentially (Endeve et al., 2012). Due to strong shear flows at the neutron star surface, magnetic fields of order $\gtrsim 10^{14}$ G may be achievable. However, the consequences of SASI-driven field amplification have not yet been investigated in self-consistent 3D CCSN simulations with neutrino transport, and the effects on the gravitational-wave emission have not yet been investigated.
@@
Another important feature of the gravitational-wave signals from CCSNe is the gravitational-wave memory. General relativity predicts that matter or radiation from an asymmetric source should result in a permanent deformation of the space-time metric, known as gravitational-wave memory. In supernova explosions, such a memory signal can arise from the asymmetric emission of neutrinos and from strongly asymmetric shock expansion (Epstein, 1978; Murphy et al., 2009; Müller et al., 2012a), and results in gravitational-wave emission below the frequency band of current gravitational-wave detectors. Exactly how low the spectrum of the memory signal reaches down in frequency is difficult to determine, as the minimum frequency that we can predict is equal to the inverse of the simulation time, and long duration simulations in 3D are too computationally expensive. However, many recent 3D studies have included the gravitational-wave memory from matter, and a few 3D studies have started to also include the low frequency gravitational-wave memory from asymmetric neutrino emission (Vartanyan et al., 2023). The neutrino memory signal is expected to be a promising source for moon-based gravitational-wave detectors (Jani & Loeb, 2020; Harms et al., 2021; Ajith et al., 2024) and space based gravitational-wave detectors such as DECIGO (Kawamura et al., 2021) and LISA (Amaro-Seoane et al., 2017). Some authors have produced analytical models to estimate the detectability of the CCSN low frequency signal in space detectors, with some studies estimating that these signals can be detected out to distances of 10 Mpc, which would result in much higher rates of CCSNe in space detectors than in current ground-based interferometers (Mukhopadhyay et al., 2021; Richardson et al., 2022; Gill, 2024).
@@
To further understand the gravitational-wave signal from CCSNe in the presence of magnetic fields and rotation, we perform four magnetohydrodynamical simulations of a $15\,M_{\odot}$ progenitor star with two different EoS. All of the models start with a weak magnetic field strength of $10^{8}$ G in the progenitor, and two of the models are rapidly rotating. We briefly discuss the hydrodynamics of the explosion and the PNS properties to aid in our description of the features of the gravitational-wave emission. All of the models undergo rapid shock revival. We find that the magnetic fields do not significantly impact the gravitational-wave amplitudes, which are $\sim 10$ cm in our non-rotating models, and $\sim 40$ cm in our rapidly rotating models. We show how the amplitude increases to as much as $\sim 200$ cm when the component of the gravitational-wave emission due to asymmetric neutrino emission is included. We show how this improves the maximum detectable distances for CCSN sources, using multiple different source angles. The low-frequency component of the gravitational-wave signal from neutrino memory is therefore an important aspect for the CCSNe science case for next-generation gravitational-wave detectors. Some previous works have questioned whether universal relations for the g/f-mode frequency are still a good fit to the actual gravitational spectrograms predicted by magnetorotational CCSN explosion models (Powell et al., 2023). We show in this work that our models are still a good fit for the universal relations when the magnetic fields are weak.
@@
Our paper is structured as follows: In Section 2 we describe the progenitor star and the setup of our simulations. Although the focus in this work is on the gravitational-wave emission, we give a brief overview of the explosion dynamics and the remnant properties in Section 3. In Section 4, we show the gravitational-wave signals produced in our simulations. In Section 5, we show the low frequency gravitational-wave emission caused by neutrino emission anisotropy. The directional dependence of the gravitational-wave emission and the prospects for the detection of the gravitational-wave signals are investigated in Section 6. In Section 7, we analyse whether the magnetic fields and rotation impact the models fit to several universal relations proposed by different groups. A discussion and conclusions follow in Section 8.
@@ 2 Progenitor model and simulation methodology
Table 1: For each model we list the model name, the equation of state, the initial core rotation rate, the initial magnetic field strength, the diagnostic explosion energies, the proto-neutron star mass at the end of the simulation time, and the gravitational-wave energy at the end of the simulation time. Rotating models reach higher explosion energies and have stronger gravitational-wave signals. Rotation also results in the formation of less massive neutron stars due to the earlier onset of the explosion.
@@
Here $\rho$, $\mathbf{v}$, $P$, and $e$, and $\mathbf{B}$ are the standard magnetohydrodynamic variables density, velocity, pressure, total gas energy density and magnetic field. Among the other variables, $c_{\mathrm{h}}$ denotes the hyperbolic cleaning speed, $\tau$ the damping time for the Lagrangian multiplier, and $Q_{\mathrm{e}}$ and $\mathbf{Q}_{\mathrm{m}}$ are the neutrino energy and momentum source terms, and $\hat{e}=e+(B^{2}+\psi^{2})/(8\pi)+$ is the total energy density of the fluid, the magnetic field and the cleaning field. The cleaning speed $c_{\mathrm{h}}$ is identified with the fast magnetosonic velocity, and the damping time is set to eight times the magnetosonic crossing time of a cell. The effective potential of Müller et al. (2008) is used to approximate the effects of relativistic gravity. The GW emission is extracted by the time-integrated quadrupole formula (Finn, 1989; Finn & Evans, 1990; Blanchet et al., 1990). Neutrinos are treated using the FMT (Fast Multi-group Transport) method of Müller & Janka (2015). Graviational redshift is included in the neutrino transport.
@@ 3 Hydrodynamic Evolution
The rapidly rotating models undergo shock revival earlier at $\sim 100$ ms after core bounce. Earlier shock revival times in magnetohydrodynamic simulations of rapidly rotating progenitors is consistent with previous work (Winteler et al., 2012; Obergaulinger et al., 2014; Mösta et al., 2014; Powell et al., 2023; Obergaulinger & Aloy, 2022). It is remarkable that such a fast explosion develops even for the low initial field strength assumed in models SFH_rr and SFHx_rr. However, a closer look at the explosions reveals that although jet-like outflows are present they are not dominated by strong jets. The evolution of the rapidly rotating models in meridional 2D slices of the entropy, at 200 ms and 400 ms post bounce, are shown in Figure 3. The SFHo_rr model has jet-like structures in both directions, however the jet is larger in the south direction. These jets do not remain collimated, but are very unstable, either because of the kink instability (Mösta et al., 2014; Bugli et al., 2021; Bugli et al., 2022; Obergaulinger & Aloy, 2021b, 2020; Powell et al., 2023) or because they are buffeted around by other outflows and downflows. The SFHx_rr model has no jet in the south direction, and forms a big jet-like structure in the North direction, which proves robust at later times, but is not strongly collimated and strongly distorted by instabilities. As the focus of this paper is on the gravitational-wave emission, we leave a more extensive study of the jet phenomenology for future work.
@@
The diagnostic explosion energies are shown in Figure 1, and the final energies at the end of the simulations are given in Table 1. For the non-rotating models, the energy is still growing rapidly at the time that the simulations were ended, which is typical for models with only $\sim 0.5$ s in duration, and they reach a final value similar to what is found in other simulations of non-rotating neutrino-driven models with a similar progenitor mass. For the rotating models, the energy grows very rapidly after shock revival but then quickly levels off after a few hundred milliseconds, as seen in previous models with stronger magnetic fields (Varma et al., 2022; Powell et al., 2023). The final energy is significantly smaller than our previous magnetorotational explosions in Powell et al. (2023), which had higher progenitor masses and stronger initial magnetic fields. The explosion energies are similar to what is typically observed in type Ib or Ic supernovae, but not in hypernovae. While purely neutrino-driven explosion models are able to reach such typical explosion energies, they can currently do this only on significantly longer time scales of seconds instead of hundreds of milliseconds (Müller et al., 2017; Bollig et al., 2021; Burrows et al., 2023). Our rapidly rotating models can be seen as exploring the potential gravitational-wave emission from supernovae with typical energies, but faster powering of the explosion than predicted by current neutrino-driven simulations.
@@
Although our simulations finish too early to obtain accurate final measurements of the PNS properties at least for the non-rotating models, we show the PNS masses, radius, spin and kicks in Figure 2. Kicks develop faster for the rapidly rotating models due to the earlier onset of the explosion, but they asymptote more quickly at fairly moderate values of $254\,\mathrm{kms}^{-1}$ and $139\,\mathrm{kms}^{-1}$ for the SFHo_rr and SFHx_rr models respectively. Final kicks for the non-rotating models cannot be determined yet, but at least for SFHo_nr, the kick may very well end up higher than for the rapidly rotating models. Variations in kick magnitude are likely determined by the stochastic variations of the explosion geometry evident from Figure 3 and 4.
All of the models asymptote quickly towards a final baryonic mass value. Rapid progenitor rotation results in a lower final PNS mass due to earlier shock revival, which is consistent with previous work, and there is not a significant difference in final PNS mass values between the different EoS. However, the PNS radius shrinks significantly faster for model SFHx_rr than for model SFHo_rr. The fast shrinking is ultimately caused by a slightly different angular momentum distribution in the PNS. In model SFHx_rr, the positive angular momentum gradient in the outer shells of the PNS is strong enough to inhibit convection, reducing the PNS convection zone to about half the width in model SFHo_rr. The rapidly rotating models maintain very similar angular momentum after the onset of shock revival until the end of the simulation. There is some spin-down of the PNS early on around shock revival, but the angular momentum remains high of order $\gtrsim 2.5\times 10^{48}\,\mathrm{erg}$, and a substantial amount of rotational energy could still be released on longer time scales. This is in contrast to the magnetorotational explosion models of Powell et al. (2023), where little rotational energy is left in the PNS already quite early during the explosion. The non-rotating models gain angular momentum over time due to stochastic accretion.

### v2

@@ 1 Introduction
Core-collapse supernovae (CCSNe) are the explosive deaths of massive stars. Once the star’s iron core reaches its effective Chandrasekhar mass, the core collapses until it reaches and overshoots nuclear saturation density. Due to the stiffening to the nuclear equation of state (EoS), the collapse is then halted, and a shock wave is launched outwards due to the rebound of the iron core. Due to energy losses by nuclear dissociation and neutrino losses, the shock quickly stalls, but still expands as an accretion shock to a radius of $\sim 150$ km. Shock revival for the majority of CCSNe is expected to be powered by neutrinos, which need to be sufficiently strong to deliver explosion energies of $\sim 10^{51}$ erg as in observed CCSNe. Explosions powered by the neutrino-driven mechanism have been simulated extensively to understand the dynamics, remnant properties, neutrinos, and the gravitational-wave (GW) emission of CCSNe (see Müller (2020); Abdikamalov et al. (2020); Burrows & Vartanyan (2021) for recent reviews). However, the majority of 3D simulations of neutrino-driven explosions do not include the effects of rotation and magnetic fields.
@@
CCSNe are a promising multi-messenger source for the current ground-based GW detectors Advanced LIGO (LIGO Scientific Collaboration et al., 2015), Advanced Virgo (Acernese et al., 2015) and KAGRA (Akutsu et al., 2021). No GWs have been detected from CCSNe in previous observing runs yet (e.g., Abbott et al., 2020; Szczepańczyk et al., 2023). Significant progress has been made in recent years predicting the GW emission for neutrino-driven CCSNe (Powell & Müller, 2019; Burrows et al., 2020; Morozova et al., 2018; Andresen et al., 2017; Kuroda et al., 2017; Mezzacappa et al., 2020; Pan et al., 2021; Radice et al., 2019). Although simulations from different groups can differ in their predictions of the GW amplitudes and precise frequency trajectories, all groups have found the dominant feature of the GW signal to be the high frequency g/f-mode. Our understanding of the GW emission modes from 3D simulations has enabled several groups to develop phenomenological models to aid CCSN GW parameter estimation and searches (Astone et al., 2018; Powell & Müller, 2022; Bruel et al., 2023).
@@
The GW signal from CCSNe with rotation and magnetic fields is not so well understood. Energetic explosions with rapid rotation may have enough GW energy to be detected outside the Milky Way (Szczepańczyk et al., 2021). Rotation results in a spike in the time-series at the core-bounce time, which has been extensively studied (Dimmelmeier et al., 2002; Richers et al., 2017; Abdikamalov et al., 2014), including the early post-bounce phase in 3D (Scheidegger et al., 2008). However, GW emission from the later phases of the explosion has not yet been investigated as thoroughly. Several studies have included rotation in their neutrino-driven explosions (Takiwaki & Kotake, 2018; Andresen et al., 2019; Powell & Müller, 2020; Shibagaki et al., 2020; Pan et al., 2021; Takiwaki et al., 2021), and a few GW signals are now available for long-duration simulations that include both magnetic fields and rotation, where the magnetic fields played a major role in driving the explosion (Jardine et al., 2022; Bugli et al., 2023; Powell et al., 2023). The f/g-mode is also found to be the dominant feature in the GW emission in simulations that include rotation and magnetic fields. However, as the effects of magnetic fields and rotation have so far been neglected in CCSN astroseismology studies, it is less clear how the GW frequency relates to the proto-neutron star (PNS) properties. One noteworthy exception is the work of Raynaud et al. (2022), which addressed later phases beyond the first $\mathord{\sim}1\,\mathrm{s}$ of evolution, and identified a clear imprint of magnetic fields from an $\alpha-\Omega$-dynamo in the PNS convection zone on GWs from inertial modes.
@@
The standing accretion shock instability (SASI) (Blondin et al., 2003; Blondin & Mezzacappa, 2006; Foglizzo et al., 2007) can also result in GW emission in the most sensitive frequency band of current GW detectors, and is often present in waveforms from neutrino-driven explosions. The GW amplitude of the SASI-driven modes is usually lower, and more difficult to detect, than the dominant g/f-mode. However, the majority of previous works investigating SASI signatures in the GW signal have not included both magnetic fields and rotation. CCSN simulations without neutrino transport have shown that SASI-driven turbulence can amplify magnetic fields exponentially (Endeve et al., 2012). Due to strong shear flows at the neutron star surface, magnetic fields of order $\gtrsim 10^{14}$ G may be achievable. However, the consequences of SASI-driven field amplification have not yet been investigated in self-consistent 3D CCSN simulations with neutrino transport, and the effects on the GW emission have not yet been investigated.
@@
Another important feature of the GW signals from CCSNe is the GW memory. General relativity predicts that matter or radiation from an asymmetric source should result in a permanent deformation of the space-time metric, known as GW memory. In supernova explosions, such a memory signal can arise from the asymmetric emission of neutrinos and from strongly asymmetric shock expansion (Epstein, 1978; Murphy et al., 2009; Müller et al., 2012a), and results in GW emission below the frequency band of current GW detectors. Exactly how low the spectrum of the memory signal reaches down in frequency is difficult to determine, as the minimum frequency that we can predict is equal to the inverse of the simulation time, and long duration simulations in 3D are too computationally expensive. However, many recent 3D studies have included the GW memory from matter, and a few 3D studies have started to also include the low frequency GW memory from asymmetric neutrino emission (Vartanyan et al., 2023). The neutrino memory signal is expected to be a promising source for moon-based GW detectors (Jani & Loeb, 2020; Harms et al., 2021; Ajith et al., 2024) and space based GW detectors such as DECIGO (Kawamura et al., 2021) and LISA (Amaro-Seoane et al., 2017). Some authors have produced analytical models to estimate the detectability of the CCSN low frequency signal in space detectors, with some studies estimating that these signals can be detected out to distances of 10 Mpc, which would result in much higher rates of CCSNe in space detectors than in current ground-based interferometers (Mukhopadhyay et al., 2021; Richardson et al., 2022; Gill, 2024).
@@
To further understand the GW signal from CCSNe in the presence of magnetic fields and rotation, we perform four magnetohydrodynamical simulations of a $15\,M_{\odot}$ progenitor star with two different EoS. All of the models start with a weak magnetic field strength of $10^{8}$ G in the progenitor, and two of the models are rapidly rotating. We briefly discuss the hydrodynamics of the explosion and the PNS properties to aid in our description of the features of the GW emission. All of the models undergo rapid shock revival. We find that the magnetic fields do not significantly impact the GW amplitudes, which are $\sim 10$ cm in our non-rotating models, and $\sim 40$ cm in our rapidly rotating models. We show how the amplitude increases to as much as $\sim 200$ cm when the component of the GW emission due to asymmetric neutrino emission is included. We show how this improves the maximum detectable distances for CCSN sources, using multiple different source angles. The low-frequency component of the GW signal from neutrino memory is therefore an important aspect for the CCSNe science case for next-generation GW detectors. Some previous works have questioned whether universal relations for the g/f-mode frequency are still a good fit to the actual gravitational spectrograms predicted by magnetorotational CCSN explosion models (Powell et al., 2023). We show in this work that our models are still a good fit for the universal relations when the magnetic fields are weak.
@@
Our paper is structured as follows: In Section 2 we describe the progenitor star and the setup of our simulations. Although the focus in this work is on the GW emission, we give a brief overview of the explosion dynamics and the remnant properties in Section 3. In Section 4, we show the GW signals produced in our simulations. In Section 5, we show the low frequency GW emission caused by neutrino emission anisotropy. The directional dependence of the GW emission and the prospects for the detection of the GW signals are investigated in Section 6. In Section 7, we analyse whether the magnetic fields and rotation impact the models fit to several universal relations proposed by different groups. A discussion and conclusions follow in Section 8.
@@ 2 Progenitor model and simulation methodology
Table 1: For each model we list the model name, the equation of state, the initial core rotation rate, the initial magnetic field strength, the diagnostic explosion energies, the proto-neutron star mass at the end of the simulation time, and the GW energy at the end of the simulation time. Rotating models reach higher explosion energies and have stronger GW signals. Rotation also results in the formation of less massive neutron stars due to the earlier onset of the explosion.
@@
Here $\rho$, $\mathbf{v}$, $P$, and $e$, and $\mathbf{B}$ are the standard magnetohydrodynamic variables density, velocity, pressure, total gas energy density and magnetic field. Among the other variables, $c_{\mathrm{h}}$ denotes the hyperbolic cleaning speed, $\tau$ the damping time for the Lagrangian multiplier, and $Q_{\mathrm{e}}$ and $\mathbf{Q}_{\mathrm{m}}$ are the neutrino energy and momentum source terms, and $\hat{e}=e+(B^{2}+\psi^{2})/(8\pi)+$ is the total energy density of the fluid, the magnetic field and the cleaning field. The cleaning speed $c_{\mathrm{h}}$ is identified with the fast magnetosonic velocity, and the damping time is set to eight times the magnetosonic crossing time of a cell. The effective potential of Müller et al. (2008) is used to approximate the effects of relativistic gravity. The GW emission is extracted by the time-integrated quadrupole formula (Finn, 1989; Finn & Evans, 1990; Blanchet et al., 1990). GW amplitudes are given as distance-normalized quadrupole amplitudes (Thorne, 1980) $hR$ (strain $h$ times distance $R$), except where the distance is explicitly specified. Neutrinos are treated using the FMT (Fast Multi-group Transport) method of Müller & Janka (2015). The GW signals are upgraded from 2D to 3D at post-bounce times of 5 ms for SFHx_rr, 3 ms for SFHo_rr, 31 ms for SFHx_nr and 13 ms for SFHo_nr. Graviational redshift is included in the neutrino transport.
@@ 3 Hydrodynamic Evolution
The rapidly rotating models undergo shock revival earlier at $\sim 100$ ms after core bounce. Earlier shock revival times in magnetohydrodynamic simulations of rapidly rotating progenitors is consistent with previous work (Winteler et al., 2012; Obergaulinger et al., 2014; Mösta et al., 2014; Powell et al., 2023; Obergaulinger & Aloy, 2022). It is remarkable that such a fast explosion develops even for the low initial field strength assumed in models SFH_rr and SFHx_rr. However, a closer look at the explosions reveals that although jet-like outflows are present they are not dominated by strong jets. The evolution of the rapidly rotating models in meridional 2D slices of the entropy, at 200 ms and 400 ms post bounce, are shown in Figure 3. The SFHo_rr model has jet-like structures in both directions, however the jet is larger in the south direction. These jets do not remain collimated, but are very unstable, either because of the kink instability (Mösta et al., 2014; Bugli et al., 2021; Bugli et al., 2022; Obergaulinger & Aloy, 2021b, 2020; Powell et al., 2023) or because they are buffeted around by other outflows and downflows. The SFHx_rr model has no jet in the south direction, and forms a big jet-like structure in the North direction, which proves robust at later times, but is not strongly collimated and strongly distorted by instabilities. As the focus of this paper is on the GW emission, we leave a more extensive study of the jet phenomenology for future work.
@@
The diagnostic explosion energies are shown in Figure 1, and the final energies at the end of the simulations are given in Table 1. For the non-rotating models, the energy is still growing rapidly at the time that the simulations were ended, which is typical for models with only $\sim 0.5$ s in duration, and they reach a final value similar to what is found in other simulations of non-rotating neutrino-driven models with a similar progenitor mass. For the rotating models, the energy grows very rapidly after shock revival but then quickly levels off after a few hundred milliseconds, as seen in previous models with stronger magnetic fields (Varma et al., 2022; Powell et al., 2023). The final energy is significantly smaller than our previous magnetorotational explosions in Powell et al. (2023), which had higher progenitor masses and stronger initial magnetic fields. The explosion energies are similar to what is typically observed in type Ib or Ic supernovae, but not in hypernovae. While purely neutrino-driven explosion models are able to reach explosion energies of $10^{51}$ erg, they can currently do this only on significantly longer time scales of seconds instead of hundreds of milliseconds (Müller et al., 2017; Bollig et al., 2021; Burrows et al., 2023). Our rapidly rotating models can be seen as exploring the potential GW emission from supernovae with typical energies, but faster powering of the explosion than predicted by current neutrino-driven simulations.
@@
Although our simulations finish too early to obtain accurate final measurements of the PNS properties at least for the non-rotating models, we show the PNS masses, radius, spin and kicks in Figure 4. Kicks develop faster for the rapidly rotating models due to the earlier onset of the explosion, but they asymptote more quickly at moderate values of $254\,\mathrm{kms}^{-1}$ and $139\,\mathrm{kms}^{-1}$ for the SFHo_rr and SFHx_rr models respectively. Final kicks for the non-rotating models cannot be determined yet, but at least for SFHo_nr, the kick may well end up higher than for the rapidly rotating models. Variations in kick magnitude are likely determined by the stochastic variations of the explosion geometry evident from Figure 3 and 2.
All of the models asymptote quickly towards a final baryonic mass value. Rapid progenitor rotation results in a lower final PNS mass due to earlier shock revival, which is consistent with previous work, and there is not a significant difference in final PNS mass values between the different EoS. However, the PNS radius shrinks significantly faster for model SFHx_rr than for model SFHo_rr. The fast shrinking is ultimately caused by a slightly different angular momentum distribution in the PNS. In model SFHx_rr, the positive angular momentum gradient in the outer shells of the PNS is strong enough to inhibit convection. This reduces the PNS convection zone to about half the width found in model SFHo_rr. It is not easily possible to establish the causal factors behind this difference because of feedback processes that may result in a bifurcation. EoS properties directly influence stability through the coefficient of the electron fraction gradient in the Ledoux criterion (see Jakobus et al., 2024, Equation 21). In addition, EoS-dependent neutrino opacities affect the evolution of the entropy and electron fraction and hence convective stability. If convective instability is sufficient to overcome stabilisation by angular momentum gradients, convective transport of angular momentum can reinforce instability and lock the PNS convection into an active state. If convective instability becomes too weak to overcome the stabilising angular momentum gradient, redistribution of angular momentum will be slower and convection may be unable to restart. The rapidly rotating models maintain very similar angular momentum after the onset of shock revival until the end of the simulation. There is some spin-down of the PNS early on around shock revival, but the angular momentum remains high of order $\gtrsim 2.5\times 10^{48}\,\mathrm{erg}$, and a substantial amount of rotational energy could still be released on longer time scales. This is in contrast to the magnetorotational explosion models of Powell et al. (2023), where little rotational energy is left in the PNS already quite early during the explosion. The non-rotating models gain angular momentum over time due to stochastic accretion.